. Calculate the following buffer solution using the Henderson-Hasselbalch equation.

1. Calculate the pH of the buffer solution that contain 0.20 M HOCl (Ka= 4.0 x 10-8) and 0.10 M HCl.


2. What is the pH of a buffer solution that contains 0.50 M CH3CO3H (Ka = 1.8 x 10-5 ) and 0.02 M

H2CO3?


1
Expert's answer
2022-03-23T16:19:02-0400

pH = pKA + log10 ([A-]/[HA]) = -log10KA + log10 ([A-]/[HA])

  1. pH = -log10(4.0*10-8) + log10(0.2/0.1) = 7.4 + 0.3 = 7.7
  2. pH = -log10(1.8*10-5) + log10(0.5/0.02) = 4.7 + 1.4 = 6.1